answersLogoWhite

0

The answer depends on drawing a 9 from WHAT!

User Avatar

Wiki User

∙ 12y ago

What else can I help you with?

Related Questions

What is probability of drawing a seven or a nine out of cards?

Probability of drawing a 7 or 9 is the probability of draw a 7 plus the probability of draw a 9 which is 4/52 + 4/52 = 8/52 = 2/13 = 0.1538.


What is the probability of drawing a 9 or a king from an deck of cards?

8 out of 32


A single card is drawn from a deck Find the probability of selecting a heart or a 9?

The probability of drawing a heart is 13 in 52 or 1 in 4. The probability of drawing a 9 is 4 in 52 or 1 in 13. The probability of drawing a heart or a 9 is (13 + 4 - 1) in 52, with the -1 required because the 9 of hearts intersects the solution set. (13 + 4 - 1) in 52 is 16 in 52, which is 4 in 13, or about 0.3077.


What is probability of drawing 10 of 52 cards?

If you draw 9 or fewer cards, the probability is 0. If you draw 10 or more card, the probability is 1.


What is the probability of drawing a 9 from a deck of cards?

4/52 or 1/13 or 0.0769.


What is the probability of drawing a black non-face card?

It is 18/52 = 9/26


There are 9 red and 6 blue marbles in a jar If you draw red marbles from the jar 7 times what is the experimental probability of drawing a red marble?

The answer is dependent on whether of not you replace the marbles in the jar. If you do, the probability of drawing a red marble is 9 in 15 or 60%, every time. If you do not replace the marbles, the probability of drawing a red marble is 2 in 8 or 25%.


Drawing 3 cards that are all 9 from a deck of cards?

The probability of drawing 3 cards, all with the value of 9, from a standard 52 card deck, is ~0.018%.


What is the probability of drawing a 9 or a jack in a deck of 52 cards?

There are four 9's and four jacks. If you can use all of them, you can use 8 out of 52 cards. The probability of drawing one of these cards is therefore 8/52 = 0.1538 or a 15.38 % chance.


What is the probability of drawing of exactly one 9 if you draw two cards without replacement from a deck of cards?

Let's call the chance of drawing a 9 on the first draw P(A). Since there are four 9s, P(A) is 4/52. Probability of not drawing a 9 is 1-(4/52). Each draw is independent so we multiply the probabilities. The probability of EXACTLY one 9 in two draws if P(A)P(1-A)=12/169 which is a about .071


What is the probability of drawing a black checker from a bag of 9 black checkers and 6 red checkers replacing it and drawing a red checker?

19


What is the probability of drawing a red 9?

Since there are 2 red nines in a 52 card deck, the probability would be 2/52 or 1/26.

Trending Questions
What is the standard deviation of the following data set rounded to the nearest tenth? What is the probability that a group of five random digits will contain at least 3 even digits? When Mrs Myles gave a test the scores were normally distributed with a mean of 72 and a standard deviation of 8 About 68 percent of her students scored between which two scores? What is the difference between a report and a summary? What is the average weight for 160cm tall boy that is 11? Explain what is meant when we say "data Vary". How does this variability affect the results of statistical analysis? What Identifies the quality of data in a database? If you draw a card from a standard deck of cards what is the probability of drawing a red card? Define detailed estimation? Can a mean have reminders? When all outcomes have the same probability of occurring the outcomes are? Can you have more than one period in a month? What is the probability of not getting a club in a deck of 52 cards? What would be the theoretical probability of rolling a die 50 times and rolling a 2? Why is a bar graph better than a pictograph? What is the difference between a type 1 and type 3 survivorship curve? What is the probability of prime numbers being chosen randomly from the number 1-49? what are the difference between activity duration estimating and estimating the effort required to perform an activity? What does statistical significance refer to? How do you find nth highest salary from table?